OpenStudy (anonymous):

Mole fraction of the solute in a 1.00 molal aqueous solution is .......?

OpenStudy (anonymous):

1 kg solvent----1 mole solute 1000g solvent--1 mole solute Since the solution is aqueous: 1000/18 moles of water---1 mole solute 55.55 moles of water combine with--1 moles solute total moles=55.55+1=56.55 moles of solute=1 mole fraction of solute=1/56.55=0.018(approx.)

OpenStudy (anonymous):

thank you..